What is the rate of heat transfer due to conduction through a single-pane glass window that is 3.0 m2 and 3.175 mm thick, when the exterior temperature is 37°C and the interior temperature is 24°C?
Answer #1

Solution) A = 3 m^2

L = 3.175 mm = 3.175×10^(-3) m

T1 = 37°C

T2 = 24°C

(Q/t) = ?

(Q/t) = ((K)(A)(dT))/(L)

Here K Is thermal conductivity

K = 0.96 W/m/°C

Here dT = difference in temperature

dT = T1 - T2 = 37 - 24 = 13°C

(Q/t) = ((0.96)(3)(13))/(3.175×10^(-3))

(Q/t) = 11792.12 W